This is a template for game-specific rules sets based on the Standard Rules. It’s easy enough to write such rules without it, but it can at least serve as a convenient checklist, so you probably want to at least glance through it.

In some areas it also includes FAQs, i.e. things that players will probably ask about if your rules about that area don’t cover them; consider saving time and sanity by including the information here.

Since there’s very little that has to be in a game-specific rules set, almost everything is in the comments of this file. GMs can go through those and turn them into actual appropriate material where necessarily. Usually you’ll be able to just erase most comment sections without putting in any material.

Oh, and remember to fill in the game name, GM team, and year, and the italicized stuff below this. (And to delete this section. …)

Rules for GAME NAME

Written and GM’d by LIST OF GMS

This game \begin{matrix}c \\ ◯ \end{matrix}YEAR MIT Assassins’ Guild & above GM team

GAME NAME uses the Standard Rules. Please read that document if you are not familiar with it.

(It is available at http://web.mit.edu/jemorris/Rules/ and (on Athena) in /mit/jemorris/Rules/.)

Experienced SR users, check the changebars and “Edition History” appendix.

You need not read all of the Combat Systems appendix; we use only the [ MA/MD/HP ; Phi ; Rock, Paper, Scissors ; Drive-By Darkwater ; Melee Darkwater ; Cinematic ] system. [Or “you don’t need the Combat Systems appendix, we use our own system.”]

This document covers only material that augments or overrides the Standard Rules. In particular, note [identify which sections are actually overridden—quite possibly none at all…].